# Determinism and Indeterminism as Model Artefacts: Toward a Model-Invariant Ontology of Physics

### v1

@@ 3.3 Quantum foundations
The Thermal Interpretation (Neumaier 2019; 2020) proposes a fully deterministic ontology, according to which expectation values of physical quantities evolve deterministically and constitute the primary ontic structure. On this view there is no fundamental collapse and no irreducible randomness; apparent stochasticity arises only at the level of coarse-grained description, when the detailed dynamics of interactions—typically involving degrees of freedom that have become effectively inaccessible through decoherence and environmental coupling in measurement-like contexts—are no longer empirically resolvable. By contrast, Barandes’ stochastic–quantum correspondence (2018) posits an ontology in which state evolution is intrinsically indeterministic: microstates undergo genuine stochastic transitions, collapse corresponds to real stochastic state updates, and the dynamics is governed by a generally non-Markovian law defined over equivalence classes of configuration-space histories. In order to reproduce quantum correlations, including entanglement, this law exhibits explicit dependence on entire past histories and departs from straightforward forward-time causal structure. By construction, both approaches reproduce the standard quantum-statistical predictions.
@@
Barandes’ choice to begin with configuration space is motivated by a desire to keep the central objects of study as close as possible to the familiar framework of classical mechanics. This is a representational move that is not clearly motivated by empirically observable facts, and the example of the qubit shows that it is not a natural starting point for all quantum systems. The central technical step of the construction is to represent a bistochastic transition matrix as the modulus square of a complex matrix, which can then be interpreted as a Kraus-type or unitary evolution operator on a Hilbert space. Hamiltonian dynamics, together with the Schrödinger, von Neumann, and Ehrenfest equations, then emerge given differentiability and a unitary lift of the dynamics—a lift that is mathematically guaranteed to exist within the framework, but only by virtue of representational choices built into the formalism. The construction has the virtue of not postulating Hilbert space structure at the outset, but the choice of lift itself requires further representational choices. How surprising the appearance of familiar quantum dynamical equations is considered to be given the setup is perhaps a matter of perspective. What the framework does succeed in demonstrating is that a stochastic completion of the empirical quantum transition structure is possible with a high degree of generality. It also illustrates clearly the way in which wave-functions are not fundamental, but derive from features that vary across representations.
The equivalence classes of configuration-space histories considered in this approach are very broad, to the point that they would in practice be impossible to rule out empirically. If a much more restricted class were selected, this could in principle be tested via multi-time correlations probed by minimally invasive POVM measurements. Given that entanglement statistics cannot be reproduced in this scheme by strictly forward-time causal dynamics alone, one natural alternative would be to restrict attention to retrocausal histories that remain local in both space and time. The purpose of pointing this out, however, is not to advocate any particular restriction, but to emphasise that a range of distinct choices—each with different conceptual advantages—is available. As the framework currently stands, the trajectories treated as ontic by the formalism are not empirically accessible at all, which raises the question of what explanatory role they play beyond illustrating one possible class of representational realisations of the invariant empirical structure.
@@
However, by insisting on the universal ontological status of this level of description, the thermal interpretation risks elevating a descriptively stable but representation-dependent structure to a status that is not empirically warranted. On Neumaier’s view, stochastic outcomes are still understood as arising from coarse-graining over strongly decohered pointer states, but the expectation values associated with these states are then identified with an operator algebra that is taken to be ontic. From the perspective adopted in this paper, this move conflates empirical stability with model invariance. Decoherence of apparatus and environment-entangled degrees of freedom can indeed be studied in a controlled and systematic way, and it is entirely plausible that the robust expectation values that emerge represent a stable endpoint of empirical refinement. What has been achieved, however, is to relocate the uncontrollable degrees of freedom to deep within the empirically inaccessible domain, where it is to be expected that alternative theoretical descriptions—potentially very different in character—could equally well mediate between the underlying dynamics and the observed, apparently stochastic outcomes.
@@ Acknowledgements
I would like to thank Nicholas Parkin for conversations, and for drawing my attention to some relevant physics work.

### v2

@@ 3.3 Quantum foundations
The Thermal Interpretation (Neumaier 2019a,b) proposes a fully deterministic ontology, according to which expectation values of physical quantities evolve deterministically and constitute the primary ontic structure. On this view there is no fundamental collapse and no irreducible randomness; apparent stochasticity arises only at the level of coarse-grained description, when the detailed dynamics of interactions—typically involving degrees of freedom that have become effectively inaccessible through decoherence and environmental coupling in measurement-like contexts—are no longer empirically resolvable. By contrast, Barandes’ stochastic–quantum correspondence (2018) posits an ontology in which state evolution is intrinsically indeterministic: microstates undergo genuine stochastic transitions, collapse corresponds to real stochastic state updates, and the dynamics is governed by a generally non-Markovian law defined over equivalence classes of configuration-space histories. In order to reproduce quantum correlations, including entanglement, this law exhibits explicit dependence on entire past histories and departs from straightforward forward-time causal structure. By construction, both approaches reproduce the standard quantum-statistical predictions.
@@
Barandes’ choice to begin with configuration space is motivated by a desire to keep the central objects of study as close as possible to the familiar framework of classical mechanics. This is a representational move that is not clearly motivated by empirically observable facts, and the example of the qubit shows that it is not a natural or privileged starting point for all quantum systems, though analogous constructions are possible in such cases. The central technical step of the construction is to represent a bistochastic transition matrix as the modulus square of a complex matrix, which can then be interpreted as a Kraus-type or unitary evolution operator on a Hilbert space. Hamiltonian dynamics, together with the Schrödinger, von Neumann, and Ehrenfest equations, then emerge given differentiability and a unitary lift of the dynamics—a lift that is mathematically guaranteed to exist within the framework, but only by virtue of representational choices built into the formalism. The construction has the virtue of not postulating Hilbert space structure at the outset, but the choice of lift itself requires further representational choices. How surprising the appearance of familiar quantum dynamical equations is considered to be given the setup is perhaps a matter of perspective. What the framework does succeed in demonstrating is that a stochastic representation of the empirical quantum transition structure is possible with a high degree of generality, without committing to any particular dynamical completion. It also illustrates clearly the way in which wave-functions are not fundamental, but derive from features that vary across representations.
The equivalence classes of configuration-space histories considered in this approach are very broad, to the point that they would in practice be impossible to rule out empirically. If a much more restricted class were selected, this could in principle be tested via multi-time correlations probed by minimally invasive POVM measurements. As the framework currently stands, however, no such restriction is imposed: the formalism is deliberately noncommittal about the structure of any intermediate dynamics, specifying only the admissible joint statistics of observable events. The purpose of noting the possibility of further restriction is therefore not to advocate any particular choice, but to emphasise that a range of distinct possibilities for the underlying dynamics—each with different conceptual advantages—remains available in principle but unresolved. In Barandes’ formulation, the configuration-space histories that appear in the formalism function merely as a stochastic form of bookkeeping for empirical correlations, and could perhaps best be seen as agnostic about the underlying dynamics beyond the empirical horizon.
@@
However, by insisting on the universal ontological status of this level of description, the thermal interpretation risks elevating a descriptively stable but representation-dependent structure to a status that is not empirically warranted. On Neumaier’s view, stochastic outcomes are still understood as arising from coarse-graining over strongly decohered pointer states, but the expectation values associated with these states are taken as ontic regardless of empirical accessibility. From the perspective adopted in this paper, this move treats invariance under unitary equivalence within a fixed quantum framework as sufficient for ontological commitment, rather than requiring invariance across all empirically admissible models (quantum or otherwise) that reproduce the same observable structure.
Decoherence of apparatus and environment-entangled degrees of freedom can indeed be studied in a controlled and systematic way, and it is entirely plausible that the robust expectation values that emerge represent a stable endpoint of empirical refinement. What has been achieved, however, is to relocate the uncontrollable degrees of freedom to deep within the empirically inaccessible domain, where it is to be expected that alternative theoretical descriptions—potentially very different in character—could equally well mediate between the underlying dynamics and the observed, apparently stochastic outcomes.
@@ Acknowledgements
I would like to thank Nicholas Parkin for conversations, and for drawing my attention to some relevant physics work, and Arnold Neumeier for comments on the first draft of this paper.